OpenStudy (anonymous):

HELP ME PLEASE!!!!!!!!!!!!! Which of these ordered pairs is the best estimate of the solution of the system y = 3x + 2 and y = -1/4 x + 1? A. (0.5, 0.5) B. (-0.3, 1.1) C. (1.2, 0.8) D. (0.8, -0.5)

OpenStudy (callisto):

y = 3x + 2 -(1) y = -1/4 x + 1 -(2) (1) - (2) 0 = 13/4 x +1 x = -1/ (13/4) = ... ?

OpenStudy (anonymous):

can you write out the steps and give answer. That way I can incorporate the steps into any other problem I might have

OpenStudy (callisto):

I have written the steps..... A more detailed one y = 3x + 2 -(1) y = -1/4 x + 1 -(2) (1) - (2) y -y = (3x + 2) - (-1/4 x + 1) 0 = 3x + 1/4x + 2 - 1 0 = 13/4 x +1 x = -1/ (13/4) = ... ?

OpenStudy (callisto):

You can get the answer by using a calculator x = -1/ (13/4) = - 4 /13 = ... ?